Grade 11Physical Chemistry

A partially dried sample contained 50 percent silica and 7 percent water. The original sample contained 12 percent of water. Find the percentage of silica in the original sample.

Abhinav Acharya
12 Years agoGrade 11

1 Answer

Sumit Majumdar
ApprovedApproved Tutor Answer12 Years ago
Dear student,
Consider 100g of the partially dried clay sample:
It contains :
7g water
50g silica
43g other things.
Total non water component: 93g

Original clay sample contained:
12g water
The total non- water components must have had mass 88g, BUT in the same proportion as above:
Mass Silica = 50/93*88 = 47.3g
Mass other things = 43/93*88 = 40.7g

% composition of original sample :
Water: 12%
Silica: 47.3%
Other things: 40.7%
Regards
Sumit